# HiCMAE: Hierarchical Contrastive Masked Autoencoder for Self-Supervised Audio-Visual Emotion Recognition

### v1

@@ 4.2.2 Results on MAFW
As for unimodal results, our method still shows strong performance when compared with state-of-the-art unimodal baselines. Specifically, when only fine-tuning the audio encoder, we achieve competitive or even better results than three cutting-edge large pre-trained speech models (i.e., Wav2Vec2 baevski2020wav2vec , HuBERT hsu2021hubert , and WavLM-Plus chen2022wavlm ), while requiring significantly fewer parameters and computational costs (e.g., 8M parameters and 1G FLOPS in HiCMAE-T versus 95M parameters and 18G FLOPs in HuBERT). For the visual modality, the proposed method achieves the best trade-off between model performance and model size. For example, HiCMAE-B outperforms the previous state-of-the-art self-supervised method MAE-DFER sun2023mae which is also pre-trained on VoxCeleb2 by +0.48% UAR and +0.53% WAR, while being 2.6$\times$ smaller and using 36% fewer FLOPs.
The results of 43 compound emotions on MAFW are shown in Table 2. This task is quite challenging due to the high difficulty of distinguishing similar compound classes (e.g., ’anger_anxiety’, ’anger_disgust’, and ’anger_disgust_anxiety’) and extremely imbalanced distribution liu2022mafw . For audio-visual modalities, when compared with state-of-the-art T-MEP zhang2023transformer , our method achieves competitive or slightly better performance in terms of UAR and WAR. When evaluating macro-averaged F1-score and AUC, our largest model shows significant improvement over the best-performing T-ESFL liu2022mafw (i.e., 12.16% versus 8.44% and 85.30% versus 74.13%). For audio modality, our method achieves consistently moderate improvement over three strong baselines. Finally, for visual modality, our method brings similar performance gains when compared with the previous best results.
@@
Wav2Vec2.0 baevski2020wav2vec A 5.27 20.38 - -
HuBERT hsu2021hubert A 5.36 20.70 - -
WavLM-Plus chen2022wavlm A 5.51 21.09 - -
@@
ResNet-18 he2016deep $\times$ V 6.18 23.83 4.89 62.92
ViT dosovitskiy2020image $\times$ V 8.62 31.76 7.46 74.90
C3D tran2015learning $\times$ V 9.51 28.12 6.73 74.54
ResNet-18+LSTM liu2022mafw $\times$ V 6.93 26.60 5.56 68.86
ViT+LSTM liu2022mafw $\times$ V 8.72 32.24 7.59 75.33
C3D+LSTM liu2022mafw $\times$ V 7.34 28.19 5.67 65.65
T-ESFL liu2022mafw $\times$ V 9.15 34.35 7.18 75.63
Former-DFER zhao2021former $\times$ V 10.21 32.07 - -
T-MEP zhang2023transformer $\times$ V 9.50 31.54 - -
@@
ResNet-18+LSTM liu2022mafw $\times$ A+V 7.85 31.03 5.95 71.08
C3D+LSTM liu2022mafw $\times$ A+V 7.45 29.88 5.76 68.13
T-ESFL liu2022mafw $\times$ A+V 9.93 34.67 8.44 74.13
T-MEP* zhang2023transformer $\times$ A+V 11.51 34.11 - -
T-MEP zhang2023transformer $\times$ A+V 13.22 36.58 - -
@@
ResNet-18+MDRE yoon2018multimodal $\times$ A+V+T 9.02 33.64 - -
AMH yoon2020attentive $\times$ A+V+T 10.24 35.35 - -
Rajan et al. rajan2022cross $\times$ A+V+T 11.09 35.33 - -
T-ESFL liu2022mafw $\times$ A+V+T 9.68 35.02 8.65 74.35
T-MEP* zhang2023transformer $\times$ A+V+T 13.25 37.69 - -
T-MEP zhang2023transformer $\times$ A+V+T 15.22 39.00 - -
@@ 4.2.3 Results on DFEW and MER-MULTI
Wav2Vec2.0 baevski2020wav2vec A 95 18 36.15 43.05
HuBERT hsu2021hubert A 95 18 35.98 43.24
WavLM-Plus chen2022wavlm A 95 18 37.78 44.64
@@
C3D tran2015learning $\times$ V 78 39 42.74 53.54
R(2+1)D-18 tran2018closer $\times$ V 33 42 42.79 53.22
3D ResNet-18 hara2018can $\times$ V 33 8 46.52 58.27
EC-STFL jiang2020dfew $\times$ V - 8 45.35 56.51
ResNet-18+LSTM zhao2021former $\times$ V - 8 51.32 63.85
ResNet-18+GRU zhao2021former $\times$ V - 8 51.68 64.02
Former-DFER zhao2021former $\times$ V 18 9 53.69 65.70
CEFLNet liu2022clip $\times$ V 13 - 51.14 65.35
EST liu2023expression $\times$ V 43 - 53.43 65.85
STT ma2022spatio $\times$ V - - 54.58 66.65
NR-DFERNet li2022nr $\times$ V - 6 54.21 68.19
DPCNet wang2022dpcnet $\times$ V 51 10 57.11 66.32
IAL li2023intensity $\times$ V 19 10 55.71 69.24
M3DFEL wang2023rethinking $\times$ V - 2 56.10 69.25
T-MEP zhang2023transformer $\times$ V 5 6 54.14 65.22
DFER-CLIP zhao2023dferclip V 153 92 59.61 71.25
SVFAP sun2023svfap V 78 44 62.83 74.27
MAE-DFER sun2023mae V 85 50 63.41 74.43
@@
ResNet-18+LSTM zhang2023transformer $\times$ A+V - - 52.41 64.32
C3D+LSTM zhang2023transformer $\times$ A+V - - 53.77 65.17
AMH yoon2020attentive $\times$ A+V - - 54.48 66.51
T-MEP* zhang2023transformer $\times$ A+V 61 22 55.06 66.30
T-MEP zhang2023transformer $\times$ A+V 61 22 57.16 68.85
@@
eGeMAPS eyben2015geneva $\times$ A - - - 17.28
VGGish hershey2017cnn $\times$ A - - - 40.76
Wav2Vec2.0 baevski2020wav2vec A 95 18 51.36 51.48
HuBERT hsu2021hubert A 95 18 50.32 52.70
WavLM-Plus chen2022wavlm A 95 18 53.43 54.16
HuBERT-CH zhang2022wenetspeech A 95 18 - 61.16
@@
ResNet-MSCeleb he2016deep $\times$ V 26 - - 40.32
ResNet-ImageNet he2016deep $\times$ V 26 - - 44.91
SENet-FER2013 hu2018squeeze $\times$ V 28 - - 56.69
ResNet-FER2013 he2016deep $\times$ V 26 - - 57.44
MANet-RAFDB zhao2021learning $\times$ V 51 - - 56.19
@@
We present the performance comparison with state-of-the-art methods on DFEW jiang2020dfew in Table 3. In audio-visual setting, our HiCMAE-S significantly outperforms the previous best-supervised method T-MEP by +5.89% UAR and +5.48% WAR, while having 25% fewer parameters and similar computational costs. In unimodal settings, although our models underperform state-of-the-art unimodal baselines, they still achieve very competitive performance. It should also be noted that our method has significantly fewer parameters and FLOPs, thus making it more suitable in resource-constrained scenarios.
The results on a Chinese dataset MER-MULTI lian2023mer are shown in Table 4. Our HiCMAE-B slightly outperforms the previous best method which utilizes MA-Net zhao2021learning (supervised on a facial expression dataset) and a powerful HuBERT-CH model (pre-trained on 10k+ hours of Chinese speech data zhang2022wenetspeech ). The audio-only result of our method has a similar performance with three self-supervised models which are also pre-trained mostly on English speech data, while it is largely inferior to the state-of-the-art HuBERT-CH. We argue that this is mainly due to the large domain gap between pre-training and fine-tuning tasks. As for video-only results, our method outperforms the previous best-supervised method.
@@ 4.3.1 Datasets
CREMA-D cao2014crema is a high-quality audio-visual dataset for studying the multimodal expression and perception of acted emotions. It consists of 7,442 video clips recorded from 91 actors. The video clips are labeled with six emotions, including happiness, sadness, anger, fear, disgust, and neutral state. Since there is no official split, we follow previous studies sun2023mae ; tran2023saaml to conduct 5-fold cross-validation in a subject-independent manner. We also conduct experiments on a subset of four emotions (i.e., happiness, sadness, anger, and neutral state) and only report the result of the last fold in this setup to align with tran2023saaml .
MSP-IMPROV busso2016msp is an acted audiovisual corpus to explore emotional behaviors during conversational dyadic interactions. The conversations are designed to elicit spontaneous emotions. The dataset contains 8,438 video clips recorded in six sessions from 12 actors. Following tran2023saaml , we only use samples of four emotion categories (i.e., anger, happiness, neutral state, and sadness) and conduct 6-fold cross-validation in a session-independent manner.
RAVDESS livingstone2018ryerson is an audio-visual dataset that includes emotional speech and song. It comprises 2,880 video clips featuring 24 professional actors, each labeled with one of eight emotions (i.e., seven basic emotions and calm). In this paper, we only use the speech part consisting of 1,440 video clips. We adopt a subject-independent 6-fold cross-validation protocol for evaluation su2020msaf ; fu2021cross ; sun2023mae .
IEMOCAP busso2008iemocap contains about 12 hours of videos from 10 subjects recorded in five sessions. In this paper, we use 5,531 samples of five classes (i.e., anger, neutral state, happiness, excitement, and sadness) and follow the common practice to merge excitement into happiness to formulate a four-emotion classification task chen2023exploring ; tseng2023av . We conduct 5-fold cross-validation in a session-independent manner.
@@ 4.3.2 Results on CREMA-D
AuxFormer goncalves2022auxformer $\times$ A - - - 58.70
LR+eGeMAPS keesing2023emotion ; eyben2015geneva A - - 52.70 -
LR+wav2vec keesing2023emotion ; baevski2020wav2vec A - - 66.50 -
Wav2Vec2.0 baevski2020wav2vec A 95 18 72.57 72.41
HuBERT hsu2021hubert A 95 18 72.72 72.57
WavLM-Plus chen2022wavlm A 95 18 73.34 73.39
@@
AuxFormer goncalves2022auxformer $\times$ V - - - 53.10
VO-LSTM ghaleb2019multimodal $\times$ V - - - 66.80
Goncalves et al. goncalves2022robust $\times$ V - - - 62.20
Lei et al. lei2023audio $\times$ V - - 64.68 64.76
SVFAP sun2023svfap V 78 44 77.31 77.37
MAE-DFER sun2023mae V 85 50 77.33 77.38
@@
EF-GRU tran2022pre $\times$ A+V - - - 57.06
LF-GRU tran2022pre $\times$ A+V - - - 58.53
TFN zadeh2017tensor $\times$ A+V - - - 63.09
MATER ghaleb2020multimodal $\times$ A+V - - - 67.20
MulT Base tran2022pre A+V 38 - - 68.87
MulT Large tran2022pre A+V 89 - - 70.22
AuxFormer goncalves2022auxformer $\times$ A+V - - - 71.70
AV-LSTM ghaleb2019multimodal $\times$ A+V - - - 72.90
AV-Gating ghaleb2019multimodal $\times$ A+V - - - 74.00
Goncalves et al. goncalves2022robust $\times$ A+V - - - 77.30
Ladder Networks goncalves2023learning $\times$ A+V - - - 80.30
@@
Attn. Pooling sadok2023vector
@@
Tran et al. tran2022pre A+V - - 83.29 83.46
AuxFormer goncalves2022auxformer $\times$ A+V - - 91.10 91.62
AV-HuBERT shi2022learning A+V 103 - - 85.47
FAV-HuBERT tran2023saaml A+V 103 - 87.34 87.61
TAPT-HuBERT tran2023saaml A+V 103 - 92.78 92.84
CTAPT-HuBERT tran2023saaml A+V 103 - 90.52 90.39
AW-HuBERT tran2023saaml A+V 103 - 93.65 93.65
@@
AV-HuBERT shi2022learning A+V 103 - - 65.27
FAV-HuBERT tran2023saaml A+V 103 - 61.05 68.35
TAPT-HuBERT tran2023saaml A+V 103 - 63.95 70.46
CTAPT-HuBERT tran2023saaml A+V 103 - 60.83 68.02
AW-HuBERT tran2023saaml A+V 103 - 65.72 71.80
@@
The performance comparison with state-of-the-art methods on CREMA-D (6-class) cao2014crema is presented in Table 5. In audio-visual setting, we first find that our method outperforms two self-supervised baselines (MulT Base and Large tran2022pre ) by substantial margins (+13% WAR). It is worth noting that they are audio-visual Transformers pre-trained on VoxCeleb2 too. However, they rely on features extracted from other models instead of the raw data as input. Thus, critical information might be lost during feature extraction, leading to significantly inferior results. VQ-MAE-AV sadok2023vector is another strong self-supervised baseline. Similar to us, it is based on masked autoencoder (but requires two-stage pre-training) and also pre-trained on VoxCeleb2. When compared with two versions of VQ-MAE-AV, our method shows large performance improvement over them and can be trained in a single stage. Specifically, the smallest HiCMAE-T surpasses the best VQ-MAE-AV-12 by +3.34% WAR while using 33% fewer parameters. With the increase in model size, HiCMAE-B pushes the performance gap to even larger (+4.49% WAR), establishing a new state-of-the-art result on this dataset. These results demonstrate the effectiveness of the three-pronged strategy to promote hierarchical feature learning in HiCMAE. Finally, our method also outperforms advanced supervised baselines (e.g., Ladder Networks goncalves2023learning ).
In unimodal setting, we observe that our method still maintains competitive performance. For example, for visual modality, HiCMAE-B is slightly inferior ($<$0.2% UAR and WAR) to state-of-the-art MAE-DFER sun2023mae , but requiring significantly fewer parameters (62%) and computational costs (36% FLOPs). For audio modality, although the performance gap between HiCMAE-B and WavLM-Plus is larger (about 2% UAR and WAR), it is still acceptable as the latter is 3$\times$ larger than the former and has 4.5$\times$ more FLOPs.
In addition to the default six emotions on this dataset, we also conduct experiments on a subset with four emotions. The audio-visual results are shown in Table 6. Among these baselines, most are self-supervised methods and the HuBERT series except for AV-HuBERT shi2022learning (i.e., FAV-HuBERT tran2023saaml , TAPT-HuBERT tran2023saaml , CTAPT-HuBERT tran2023saaml , and AW-HuBERT tran2023saaml ) are pre-trained on VoxCeleb2 with industry-level computation resources (32 Tesla-V100 GPUs) for approximately 10 days. When compared with the best-performing AW-HuBERT, our HiCMAE-B still shows slight improvement (+0.35% UAR and +0.48% WAR), while being 21% smaller and training-friendly (we only need 4 Tesla-V100 GPUs to pre-train the model for about 5 days). It should also be noted that AW-HuBERT is a semi-supervised method that requires another labeled dataset and unlabeled samples from VoxCeleb2 for affective adaptation to obtain improved performance. Thus, these results amply demonstrate the superiority of the proposed method.

### v2

@@ 4.2.2 Results on MAFW
As for unimodal results, our method still shows strong performance when compared with state-of-the-art unimodal baselines. Specifically, when only fine-tuning the audio encoder, we achieve competitive or even better results than three cutting-edge large pre-trained speech models (i.e., Wav2Vec2 [47], HuBERT [48], and WavLM-Plus [71]), while requiring significantly fewer parameters and computational costs (e.g., 8M parameters and 1G FLOPS in HiCMAE-T versus 95M parameters and 18G FLOPs in HuBERT). For the visual modality, the proposed method achieves the best trade-off between model performance and model size. For example, HiCMAE-B outperforms the previous state-of-the-art self-supervised method MAE-DFER [50] which is also pre-trained on VoxCeleb2 by +0.48% UAR and +0.53% WAR, while being 2.6$\times$ smaller and using 36% fewer FLOPs.
The results of 43 compound emotions on MAFW are shown in Table 3. This task is quite challenging due to the high difficulty of distinguishing similar compound classes (e.g., ‘anger_anxiety’, ‘anger_disgust’, and ‘anger_disgust_anxiety’) and extremely imbalanced distribution [5]. For audio-visual modalities, when compared with state-of-the-art T-MEP [9], our method achieves competitive or slightly better performance in terms of UAR and WAR. When evaluating macro-averaged F1-score and AUC, our largest model shows significant improvement over the best-performing T-ESFL [5] (i.e., 12.16% versus 8.44% and 85.30% versus 74.13%). For audio modality, our method achieves consistently moderate improvement over three strong baselines. Finally, for visual modality, our method brings similar performance gains when compared with the previous best results.
@@
Wav2Vec2.0 [47] A 5.27 20.38 - -
HuBERT [48] A 5.36 20.70 - -
WavLM-Plus [71] A 5.51 21.09 - -
@@
ResNet-18 [72] $\times$ V 6.18 23.83 4.89 62.92
ViT [73] $\times$ V 8.62 31.76 7.46 74.90
C3D [43] $\times$ V 9.51 28.12 6.73 74.54
ResNet-18+LSTM [5] $\times$ V 6.93 26.60 5.56 68.86
ViT+LSTM [5] $\times$ V 8.72 32.24 7.59 75.33
C3D+LSTM [5] $\times$ V 7.34 28.19 5.67 65.65
T-ESFL [5] $\times$ V 9.15 34.35 7.18 75.63
Former-DFER [74] $\times$ V 10.21 32.07 - -
T-MEP [9] $\times$ V 9.50 31.54 - -
@@
ResNet-18+LSTM [5] $\times$ A+V 7.85 31.03 5.95 71.08
C3D+LSTM [5] $\times$ A+V 7.45 29.88 5.76 68.13
T-ESFL [5] $\times$ A+V 9.93 34.67 8.44 74.13
T-MEP* [9] $\times$ A+V 11.51 34.11 - -
T-MEP [9] $\times$ A+V 13.22 36.58 - -
@@
ResNet-18+MDRE [77] $\times$ A+V+T 9.02 33.64 - -
AMH [76] $\times$ A+V+T 10.24 35.35 - -
Rajan et al. [78] $\times$ A+V+T 11.09 35.33 - -
T-ESFL [5] $\times$ A+V+T 9.68 35.02 8.65 74.35
T-MEP* [9] $\times$ A+V+T 13.25 37.69 - -
T-MEP [9] $\times$ A+V+T 15.22 39.00 - -
@@ 4.2.3 Results on DFEW and MER-MULTI
Wav2Vec2.0 [47] A 95 18 36.15 43.05
HuBERT [48] A 95 18 35.98 43.24
WavLM-Plus [71] A 95 18 37.78 44.64
@@
C3D [43] $\times$ V 78 39 42.74 53.54
R(2+1)D-18 [82] $\times$ V 33 42 42.79 53.22
3D ResNet-18 [83] $\times$ V 33 8 46.52 58.27
EC-STFL [29] $\times$ V - 8 45.35 56.51
ResNet-18+LSTM [74] $\times$ V - 8 51.32 63.85
ResNet-18+GRU [74] $\times$ V - 8 51.68 64.02
Former-DFER [74] $\times$ V 18 9 53.69 65.70
CEFLNet [84] $\times$ V 13 - 51.14 65.35
EST [85] $\times$ V 43 - 53.43 65.85
STT [86] $\times$ V - - 54.58 66.65
NR-DFERNet [87] $\times$ V - 6 54.21 68.19
DPCNet [88] $\times$ V 51 10 57.11 66.32
IAL [89] $\times$ V 19 10 55.71 69.24
M3DFEL [90] $\times$ V - 2 56.10 69.25
T-MEP [9] $\times$ V 5 6 54.14 65.22
DFER-CLIP [75] V 153 92 59.61 71.25
SVFAP [49] V 78 44 62.83 74.27
MAE-DFER [50] V 85 50 63.41 74.43
@@
ResNet-18+LSTM [9] $\times$ A+V - - 52.41 64.32
C3D+LSTM [9] $\times$ A+V - - 53.77 65.17
AMH [76] $\times$ A+V - - 54.48 66.51
T-MEP* [9] $\times$ A+V 61 22 55.06 66.30
T-MEP [9] $\times$ A+V 61 22 57.16 68.85
@@
eGeMAPS [32] $\times$ A - - - 17.28
VGGish [41] $\times$ A - - - 40.76
Wav2Vec2.0 [47] A 95 18 51.36 51.48
HuBERT [48] A 95 18 50.32 52.70
WavLM-Plus [71] A 95 18 53.43 54.16
HuBERT-CH [91] A 95 18 - 61.16
@@
ResNet-MSCeleb [72] $\times$ V 26 - - 40.32
ResNet-ImageNet [72] $\times$ V 26 - - 44.91
SENet-FER2013 [92] $\times$ V 28 - - 56.69
ResNet-FER2013 [72] $\times$ V 26 - - 57.44
MANet-RAFDB [93] $\times$ V 51 - - 56.19
@@
We present the performance comparison with state-of-the-art methods on DFEW [29] in Table 4. In the audio-visual setting, our HiCMAE-S significantly outperforms the previous best-supervised method T-MEP by +5.89% UAR and +5.48% WAR, while having 25% fewer parameters and similar computational costs. In unimodal settings, although our models underperform state-of-the-art unimodal baselines, they still achieve very competitive performance. It should also be noted that our method has significantly fewer parameters and FLOPs, thus making it more suitable in resource-constrained scenarios.
The results on a Chinese dataset MER-MULTI [25] are shown in Table 5. Our HiCMAE-B slightly outperforms the previous best method which utilizes MA-Net [93] (supervised on a facial expression dataset) and a powerful HuBERT-CH model (pre-trained on 10k+ hours of Chinese speech data [91]). The audio-only result of our method has a similar performance with three self-supervised models which are also pre-trained mostly on English speech data, while it is largely inferior to the state-of-the-art HuBERT-CH. We argue that this is mainly due to the large domain gap between pre-training and fine-tuning tasks. As for video-only results, our method outperforms the previous best-supervised method.
@@ 4.3.1 Datasets
CREMA-D [28] is a high-quality audio-visual dataset for studying the multimodal expression and perception of acted emotions. It consists of 7,442 video clips recorded by 91 actors. The video clips are labeled with six emotions, including happiness, sadness, anger, fear, disgust, and neutral state. Since there is no official split, we follow previous studies [50, 94] to conduct 5-fold cross-validation in a subject-independent manner. We also conduct experiments on a subset of four emotions (i.e., happiness, sadness, anger, and neutral state) and only report the result of the last fold in this setup to align with [94].
MSP-IMPROV [68] is an acted audiovisual corpus to explore emotional behaviors during conversational dyadic interactions. The conversations are designed to elicit spontaneous emotions. The dataset contains 8,438 video clips recorded in six sessions from 12 actors. Following [94], we only use samples of four emotion categories (i.e., anger, happiness, neutral state, and sadness) and conduct 6-fold cross-validation in a session-independent manner.
RAVDESS [69] is an audio-visual dataset that includes emotional speech and song. It comprises 2,880 video clips featuring 24 professional actors, each labeled with one of eight emotions (i.e., seven basic emotions and calm). In this paper, we only use the speech part consisting of 1,440 video clips. We adopt a subject-independent 6-fold cross-validation protocol for evaluation [95, 96, 50].
IEMOCAP [70] contains about 12 hours of videos from 10 subjects recorded in five sessions. In this paper, we use 5,531 samples of five classes (i.e., anger, neutral state, happiness, excitement, and sadness) and follow the common practice of merging excitement into happiness to formulate a four-emotion classification task [97, 98]. We conduct 5-fold cross-validation in a session-independent manner.
@@ 4.3.2 Results on CREMA-D
AuxFormer [53] $\times$ A - - - 58.70
LR+eGeMAPS [99, 32] A - - 52.70 -
LR+wav2vec [99, 47] A - - 66.50 -
Wav2Vec2.0 [47] A 95 18 72.57 72.41
HuBERT [48] A 95 18 72.72 72.57
WavLM-Plus [71] A 95 18 73.34 73.39
@@
AuxFormer [53] $\times$ V - - - 53.10
VO-LSTM [100] $\times$ V - - - 66.80
Goncalves et al. [101] $\times$ V - - - 62.20
Lei et al. [102] $\times$ V - - 64.68 64.76
SVFAP [49] V 78 44 77.31 77.37
MAE-DFER [50] V 85 50 77.33 77.38
@@
EF-GRU [55] $\times$ A+V - - - 57.06
LF-GRU [55] $\times$ A+V - - - 58.53
TFN [103] $\times$ A+V - - - 63.09
MATER [104] $\times$ A+V - - - 67.20
MulT Base [55] A+V 38 - - 68.87
MulT Large [55] A+V 89 - - 70.22
AuxFormer [53] $\times$ A+V - - - 71.70
AV-LSTM [100] $\times$ A+V - - - 72.90
AV-Gating [100] $\times$ A+V - - - 74.00
Goncalves et al. [101] $\times$ A+V - - - 77.30
Ladder Networks [105] $\times$ A+V - - - 80.30
@@
Attn. Pooling [20]
@@
Tran et al. [55] A+V - - 83.29 83.46
AuxFormer [53] $\times$ A+V - - 91.10 91.62
AV-HuBERT [106] A+V 103 - - 85.47
FAV-HuBERT [94] A+V 103 - 87.34 87.61
TAPT-HuBERT [94] A+V 103 - 92.78 92.84
CTAPT-HuBERT [94] A+V 103 - 90.52 90.39
AW-HuBERT [94] A+V 103 - 93.65 93.65
@@
AV-HuBERT [106] A+V 103 - - 65.27
FAV-HuBERT [94] A+V 103 - 61.05 68.35
TAPT-HuBERT [94] A+V 103 - 63.95 70.46
CTAPT-HuBERT [94] A+V 103 - 60.83 68.02
AW-HuBERT [94] A+V 103 - 65.72 71.80
@@
The performance comparison with state-of-the-art methods on CREMA-D (6-class) [28] is presented in Table 6. In the audio-visual setting, we first find that our method outperforms two self-supervised baselines (MulT Base and Large [55]) by substantial margins (+13% WAR). It is worth noting that they are audio-visual Transformers pre-trained on VoxCeleb2 too. However, they rely on features extracted from other models instead of the raw data as input. Thus, critical information might be lost during feature extraction, leading to significantly inferior results. VQ-MAE-AV [20] is another strong self-supervised baseline. Similar to us, it is based on masked autoencoder (but requires two-stage pre-training) and also pre-trained on VoxCeleb2. When compared with two versions of VQ-MAE-AV, our method shows large performance improvement over them and can be trained in a single stage. Specifically, the smallest HiCMAE-T surpasses the best VQ-MAE-AV by +3.34% WAR while using 33% fewer parameters. With the increase in model size, HiCMAE-B pushes the performance gap to even larger (+4.49% WAR), establishing a new state-of-the-art result on this dataset. These results demonstrate the effectiveness of the three-pronged strategy to promote hierarchical feature learning in HiCMAE. Finally, our method also outperforms advanced supervised baselines (e.g., Ladder Networks [105]).
In the unimodal setting, we observe that our method still maintains competitive performance. For example, for visual modality, HiCMAE-B is slightly inferior ($<$0.2% UAR and WAR) to state-of-the-art MAE-DFER [50], but requires significantly fewer parameters (62%) and computational costs (36% FLOPs). For audio modality, although the performance gap between HiCMAE-B and WavLM-Plus is larger (about 2% UAR and WAR), it is still acceptable as the latter is 3$\times$ larger than the former and has 4.5$\times$ more FLOPs.
In addition to the default six emotions on this dataset, we also conduct experiments on a subset with four emotions. The audio-visual results are shown in Table 7. Among these baselines, most are self-supervised methods and the HuBERT series except for AV-HuBERT [106] (i.e., FAV-HuBERT [94], TAPT-HuBERT [94], CTAPT-HuBERT [94], and AW-HuBERT [94]) are pre-trained on VoxCeleb2 with industry-level computation resources (32 Tesla-V100 GPUs) for approximately 10 days. When compared with the best-performing AW-HuBERT, our HiCMAE-B still shows slight improvement (+0.35% UAR and +0.48% WAR), while being 21% smaller and training-friendly (we only need 4 Tesla-V100 GPUs to pre-train the model for about 5 days). It should also be noted that AW-HuBERT is a semi-supervised method that requires another labeled dataset and unlabeled samples from VoxCeleb2 for affective adaptation to obtain improved performance. Thus, these results amply demonstrate the superiority of the proposed method.